OpenStudy (anonymous):

Does anyone know how to find Q1,IRQ,Q3??

OpenStudy (anonymous):

Q1 = 25% or pop IQR = Q3-Q1 = 50% Q3 = 75%

OpenStudy (anonymous):

My data is 30,35,20,25,42,32,3,20,35

OpenStudy (anonymous):

how do i do that

OpenStudy (anonymous):

\[Q1 = \frac{1}{4}(n+1)\text{ Where n is the number of data values }\] so we have n = 9 \[\frac{1}{4}(9+1) = \text{2.5 th value}\] so take the average of the second and third values {3, 20, 20, 25, 30, 32, 35, 35, 42} \[Q1 = \frac{20+20}{2}=20\]

OpenStudy (anonymous):

\[Q_3 = \frac{3}{4}(n+1)\] \[Q_3 = \frac{3}{4}(10) = 7.5 \text{ th value }\] so take the average of the 7th and 8th values of {3, 20, 20, 25, 30, 32, 35, 35, 42} \[Q_3 = \frac{35+35}{2}=35\] The interquatile Range \[ IQR = Q_3-Q_1 = 35-20 = 15 \]

OpenStudy (anonymous):

Or you could do 7.5-2.5 = 5 so your 5th value in {3, 20, 20, 25, 30, 32, 35, 35, 42} divided by 2 = 30/2 = 15

OpenStudy (anonymous):

\[\text{IQR = } \text{Q}_3-\text{Q}_2 = 35-20 = 15 \]